# An Invitation to Neuroalgebraic Geometry

### v1

@@ Abstract
In this expository work, we promote the study of function spaces parameterized by machine learning models through the lens of algebraic geometry. To this end, we focus on algebraic models, such as neural networks with polynomial activations, whose associated function spaces are semi-algebraic varieties. We outline a dictionary between algebro-geometric invariants of these varieties, such as dimension, degree, and singularities, and fundamental aspects of machine learning, such as sample complexity, expressivity, training dynamics, and implicit bias. Along the way, we review the literature and discuss ideas beyond the algebraic domain. This work lays the foundations of a research direction bridging algebraic geometry and deep learning, that we refer to as neuroalgebraic geometry.
@@ 1.1 The Power of Algebraic Geometry
A rich class of machine learning models are (semi)-algebraic, meaning that the corresponding function is a (piece-wise) polynomial in both the input and the parameters. Examples include deep neural networks with polynomial or Rectified Linear Unit (ReLU) activation function.
For (semi-) algebraic models, the neuromanifold can be defined by a finite set of polynomial equalities and inequalities, i.e., it is a semi-algebraic variety. The study of these spaces is central to algebraic geometry, a field with a rich mathematical history that offers a wealth of ideas, tools, and invariants. Many of these, as we will argue, are relevant for machine learning, and can shed light on important questions in the field. For example, a fundamental invariant of algebraic varieties is their degree, which is characteristic of algebraic geometry since it is undefined for, say, general differentiable manifolds. The degree plays a central role in learning aspects of the corresponding model (see Section 3.1). Another crucial advantage of algebraic geometry is that, unlike differential geometry, it naturally encompasses singular spaces. This is convenient in the context of deep learning; despite the ‘manifold’ terminology, neuromanifolds of neural networks are far from being smooth spaces, since they often exhibit singularities that affect the behavior of the corresponding model. Moreover, as previously mentioned, distances in the ambient space of neuromanifolds play a central role in the learning process of machine learning models. This suggests that the metric structure should be taken into account when analyzing neuromanifolds. Distance-based problems over algebraic varieties are the focus of the field of metric algebraic geometry (Breiding et al., 2024), from which we will borrow several tools throughout this work.
@@ 1.2 Neuroalgebraic Geometry
Based on these reflections, we promote the study of machine learning models via the powerful toolbox of algebraic geometry, potentially leading to new insights in deep learning theory. We propose the term neuroalgebraic geometry to refer to this emerging field of research. Neuroalgebraic geometry is closely related to algebraic statistics (Pistone et al., 2000) – the study of statistical models defined by polynomial equations – and can be interpreted as its counterpart in the context of machine learning.
@@ 1.3 Related Approaches
Several alternative approaches have been explored to analyze deep learning models, and in particular the function spaces they parameterize. Each of these approaches advocates for a specific mathematical formalism, which naturally brings original ideas to the field of machine learning, but also comes with restrictions on the models it considers.
Information geometry (Amari, 2016) is a well-established field which studies statistical models and neuromanifolds from the perspective of Riemannian geometry. While similar in spirit to neuroalgebraic geometry — both study the geometry of neuromanifolds — the two approaches rely on different mathematical machinery. The Riemannian approach of information geometry is not constrained to algebraic models, but it fails to capture some fundamental aspects of neural models. As mentioned in Section 1, algebraic geometry provides richer invariants than differential and Riemannian geometry – such as the degree – and it is more suited for analyzing singular neuromanifolds. The importance of singularities for these models has led to the development of singular learning theory (Watanabe, 2009). The latter stems from information geometry, and naturally employs ideas from algebraic geometry. Neuroalgebraic geometry extends this diction, aiming to explore the full potential of a bridge between algebraic geometry and machine learning.
@@
Further examples of approaches bridging pure mathematics – in particular, algebra and geometry – with machine learning include geometric deep learning (Bronstein et al., 2021) – which is concerned with symmetry properties of neural networks, such as invariance and equivariance – and categorical deep learning (Gavranović et al., 2024) – which proposes category theory for a compositional understanding of neural architectures.
@@ 2 Preliminaries
In this section, we review the basics of machine learning and introduce the core object of study of neuroalgebraic geometry, i.e., neuromanifolds.
@@
where $\mathcal{\ell}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a loss function. When $\mathcal{Y}$ is a Euclidean space or a simplex, respectively, common choices for $\ell$ are the quadratic loss $\ell(y,\hat{y})=\|y-\hat{y}\|^{2}$ or the cross-entropy loss $\ell(y,\hat{y})=-\sum_{j}y_{j}\log\hat{y}_{j}$. The goal of ERM is to find parameters that generalize to unseen examples. This is typically formulated by assuming that datapoints are drawn from some distribution, and that the empirical objective $L_{\mathcal{D}}$ approximates $\mathbb{E}_{x,y}[\ell(y,f_{w}(x))]$, referred to as generalization error. Often the data distribution is further assumed to take the form $x\sim\pi(x)$, $y=f^{*}(x)$, where $\pi$ is a distribution of inputs and $f^{*}$ is a deterministic function mapping inputs to outputs, interpreted as the ground-truth function underlying the given task.
In deep learning, elaborated parametric models are constructed by composing simpler ones, referred to as layers. The connectivity of such layers determines an architecture class – a family of models with similar design, e.g., a Multi-Layer Perceptron (MLP) is a sequential composition
@@ 2.1 Neuromanifolds
In other words, the neuromanifold is the image of the parametrization map, denoted by $\varphi\colon\mathcal{W}\rightarrow\mathcal{M}$, $w\mapsto f_{w}$. Describing neuromanifolds is challenging, as even simple architecture classes parametrize complex families of functions. Indeed, the parametrization map is typically non-linear and not one-to-one, resulting in different parameters that identify the same function. Therefore, the neuromanifold carries intrinsic geometric structure that differs drastically from the one of the parameter space $\mathcal{W}$. As we shall argue, the geometry of $\mathcal{M}$ is related to several fundamental aspects in machine learning.
@@
The main class of examples of algebraic models is provided by neural networks with polynomial activation functions. Indeed, if $\sigma$ is a polynomial, then the expression in Equation 2 is polynomial in both $x$ and $w$. The study of MLPs with monomial activation function $\sigma(z)=z^{k}$ has been initiated by Kileel et al. (2019). For shallow networks – i.e., with $L=2$ layers – the corresponding neuromanifold coincides with the space of symmetric tensors of bounded (Waring) rank (Arjevani et al., 2025). For linear networks – i.e., with no activation function or, equivalently, with $k=1$ – the neuromanifold is, similarly, a space of matrices with bounded rank (Trager et al., 2019). This space is referred to as determinantal variety and it is well-understood – see Section A for a detailed discussion. For Convolutional Neural Networks (CNNs), the geometry of the neuromanifold differs drastically. In this case, up to rescaling the parameters, $\mathcal{M}$ is linearly birational – i.e., related by a linear map that is an isomorphism almost everywhere – to a more complex space known as the Segre-Veronese variety (Kohn et al., 2022, 2024; Shahverdi, 2024; Shahverdi et al., 2024). A further example of an algebraic model is provided by linear (or ‘lightning’) attention mechanisms. These are non-normalized versions of the layers of a Transformer (Vaswani, 2017), which are popular nowadays across several applications. Linear attention mechanisms are cubical layers, and their neuromanifolds behave similarly to CNNs (Henry et al., 2024).
@@ 2.2 Learning on the Neuromanifold
While optimizing over the parameter space enables to deploy algorithms such as gradient descent due to its Euclidean nature, over the neuromanifold the optimization problem is often more tractable from a mathematical perspective, and can be interpreted geometrically. For instance, for the quadratic loss, the ERM consists in minimizing a potentially degenerate quadratic form over $\mathcal{M}$, while the corresponding generalization error coincides with the squared $L^{2}$ distance between $f\in\mathcal{M}$ and the ground-truth function $f^{*}$. In both cases, learning amounts to solving a nearest-point problem over the neuromanifold w.r.t. the (squared) distance associated to the quadratic form.
The loss landscape is the graph of the loss $L_{\mathcal{D}}$ in parameter space. For algebraic loss functions $\mathcal{L}_{\mathcal{D}}$ (e.g., quadratic loss or Wasserstein distance for discrete distributions), the loss landscape is also a semi-algebraic variety, analogously to the neuromanifold. For the cross-entropy loss or other loss functions with algebraic derivatives, at least the most important points of the loss landscape – namely, the critical points – depend algebraically on the data $\mathcal{D}$. In summary, in these settings, training is an algebraic optimization problem. As anticipated in Section 1, these problems lie at the heart of metric algebraic geometry, whose tools, as we shall see, are central to neuroalgebraic geometry.
@@ 3.1 Dimension, Degree, and Covering Number
where $m$ and $d$ are the dimension and the degree of $\mathcal{M}$, respectively, and $C$ is a constant depending on the dimension of the ambient space $\mathcal{V}$ but independent of $m$, $d$, and $\varepsilon$.
@@ 3.2 Singularities
Interestingly, singularities of neuromanifolds frequently correspond to subnetworks of the original architecture, i.e., functions that can be represented by a smaller network from the same model class. This type of property has been observed for different architectures such as MLPs – with (Arjevani et al., 2025) and without (Trager et al., 2019) activation function – and for CNNs with monomial activation functions (Shahverdi et al., 2024). In these settings, singularities may be responsible for a form of ‘automatic model selection’, encouraging simpler models over more complex ones. This is also consistent with empirical observations in the literature, according to which the performance of a model can often be matched by sparse subnetworks obtained through pruning or distillation (Zhu et al., 2023) or with careful initialization (Frankle & Carbin, 2018).
@@ 3.3 Parameterization and Fibers
The fibers of a map $\varphi\colon\mathcal{W}\to\mathcal{M}$ describe domain points that are collapsed to the same output. Formally, the fiber of $f\in\mathcal{M}$ is $\varphi^{-1}(f)=\{w\in\mathcal{W}\ |\ \varphi(w)=f_{w}=f\}$. For an algebraic map $\varphi$, almost all its fibers are isomorphic (over $\mathbb{C}$), and collectively referred to as the generic fiber. Computing the fibers of a network’s parameterization formalizes the problem of identifiability in the context of machine learning, which is concerned with recovering the parameters of a model from the function it defines.
@@
As an example, we now describe the fibers of MLPs. Other architectures have been considered, e.g. CNNs (Shahverdi et al., 2024) and attention mechanisms (Henry et al., 2024). First, arbitrary MLPs exhibit combinatorial symmetries in their parameterization, by permuting the output of a layer and permuting back the input of the following one. Formally, given a permutation matrix $T$ for some layer $1\leq i<L$, the simultaneous transformations $W_{i}\mapsto TW_{i}$ and $W_{i+1}\mapsto W_{i+1}T^{-1}$ do not alter the function defined by the MLP. Moreover, for (positively) homogeneous activation functions $\sigma$ – e.g., $\sigma(x)=x^{r}$ or ReLU – a similar symmetry arises by rescaling the input and output of each neuron. Lastly, when $\sigma(x)=x$ is the identity, the same holds for arbitrary invertible matrices $T$. Intuitively, non-trivial activation functions ‘break’ symmetries, reducing the latter to neuron-wise transformations, i.e., permutations and possibly rescalings. It has been conjectured that such symmetries characterize the generic fibers of algebraic MLPs (Kileel et al., 2019; Kubjas et al., 2024), which has been recently proven for monomial activations of high degree $r\gg 0$ (Finkel et al., 2024). In the non-algebraic case, the same result is well-known when $\sigma$ is the hyperbolic tangent (Fefferman et al., 1994), while for ReLU additional non-trivial symmetries arise (Grigsby et al., 2023).
@@ 3.4 Critical Points and Gradient Descent
In general, a parameter $w\in\mathrm{Crit}(\varphi)$ can be a critical point of the loss $L_{\mathcal{D}}$ even though the corresponding function $f_{w}\in\mathcal{M}$ is not a critical point for the loss functional $\mathcal{L}_{\mathcal{D}}$ in function space (see Section A for an example). These parameters are sometimes called spurious critical points (Trager et al., 2019). For some network architectures, spurious critical points will be typically avoided by gradient descent, while for others they are found with positive probability (Kohn et al., 2022). For instance, in the case of monomial activation, MLPs can have spurious critical points as local minima with positive probability (Kileel et al., 2019), while for CNNs they only correspond to the zero function $0\in\mathcal{M}$ (Shahverdi et al., 2024). For attention networks, the characterization of spurious critical points is an open problem.
@@
Figure 4: Illustration of a curve with the corresponding data discriminant (dashed) and Voronoi cell of the singularity (gray). For points inside (resp. outside) the discriminant, the Euclidean distance from the curve 2 (resp. 4) real critical points.
While critical points describe the equilibria of the gradient flow, its dynamical behavior can be described by its preserved quantities, referred to as dynamical invariants. These are relations in the parameters $w\in\mathcal{W}$ that stay constant along the curve that gradient flow traces in $\mathcal{W}$. Dynamical invariants have been discussed for a variety of network architectures (Kohn et al., 2022; Williams et al., 2019; Du et al., 2018), and the algebraic one have been recently described in detail (Marcotte et al., 2024). Knowing the invariants not only enables to analyze the training dynamics, but can be exploited to design well-behaved parameter initializations, since that gradient flow will avoid all parameters that do not satisfy the same invariants as the initial values (Arora et al., 2018). For instance, a balanced initialization can lead to stable learning dynamics.

### v2

# Algebra Unveils Deep Learning An Invitation to Neuroalgebraic Geometry

@@ Abstract
In this position paper, we promote the study of function spaces parameterized by machine learning models through the lens of algebraic geometry. To this end, we focus on algebraic models, such as neural networks with polynomial activations, whose associated function spaces are semi-algebraic varieties. We outline a dictionary between algebro-geometric invariants of these varieties, such as dimension, degree, and singularities, and fundamental aspects of machine learning, such as sample complexity, expressivity, training dynamics, and implicit bias. Along the way, we review the literature and discuss ideas beyond the algebraic domain. This work lays the foundations of a research direction bridging algebraic geometry and deep learning, that we refer to as neuroalgebraic geometry.
@@ 1.1 The Power of Algebraic Geometry
A rich class of machine learning models are (semi)-algebraic, meaning that the corresponding function is (piece-wise) polynomial in both the input and the parameters. Examples include deep neural networks with polynomial or Rectified Linear Unit (ReLU) activation function.
For (semi-) algebraic models, the neuromanifold can be defined by a finite set of polynomial equalities and inequalities, i.e., it is a semi-algebraic variety. The study of these spaces is central to algebraic geometry, a field with a rich mathematical history that offers a wealth of ideas, tools, and invariants. Many of these, as we will argue, can be interpreted from the perspective of machine learning, shedding light on important questions in the field. For example, a fundamental invariant of algebraic varieties is their degree, which is characteristic of algebraic geometry since it is undefined for, say, general differentiable manifolds. The degree plays a central role in learning aspects of the corresponding model (see Section 4.1). Another crucial advantage of algebraic geometry is that, unlike differential geometry, it naturally encompasses singular spaces. This is convenient in the context of deep learning; despite the ‘manifold’ terminology, neuromanifolds of neural networks are far from being smooth spaces, since they often exhibit singularities that affect the behavior of the corresponding model. Moreover, as previously mentioned, distances in the ambient space of neuromanifolds play a central role in the learning process of machine learning models. This suggests that the metric structure should be taken into account when analyzing neuromanifolds. Distance-based problems over algebraic varieties are the focus of the field of metric algebraic geometry (Breiding et al., 2024), from which we will borrow several tools throughout this work.
@@ 1.2 Neuroalgebraic Geometry
Based on these reflections, we argue that algebraic machine learning models can be studied using the powerful toolbox of algebraic geometry, leading to new insights in deep learning theory. We propose the term neuroalgebraic geometry to refer to this emerging field of research. Neuroalgebraic geometry is closely related to algebraic statistics (Pistone et al., 2000) – the study of statistical models defined by polynomial equations – and can be interpreted as its counterpart in the context of machine learning.
@@
2 Alternative Views
A potential counterargument to our position is that the relvance of algebraic models might be limited, both theoretically and in practice. Indeed, polynomials are rarely used as activation functions in real-life neural networks, and they are excluded by standard formulations of the universal approximation theorem for deep learning (Pinkus, 1999). However, we believe that algebraic models provide a setting where many intuitions that hold for networks in general can be formulated rigorously and, potentially, proved. Moreover, algebraic models are widely used as building blocks in actual architectures, as argued in Section 3. Regarding approximation results, we note that algebraic models can approximate arbitrary network architectures, and indeed universal approximation theorems can be extended to include polynomials by considering networks with varying depth (Yu et al., 2021).
From a broader perspective, several alternative approaches have been explored to analyze deep learning models, and in particular the function spaces they parameterize. Each of these approaches advocates for a specific mathematical formalism, which naturally brings original ideas to the field of machine learning, but also comes with restrictions on the models it considers. We believe that algebraic geometry provides unique tools to analyze aspects that are overlooked by other formalisms, as we will discuss in the next section.
@@ 2.1 Related Approaches
Information geometry (Amari, 2016) is a well-established field which studies statistical models and neuromanifolds from the perspective of Riemannian geometry. While similar in spirit to neuroalgebraic geometry — both study the geometry of neuromanifolds — the two approaches rely on different mathematical machinery. The Riemannian approach of information geometry is not constrained to algebraic models, but it fails to capture some fundamental aspects of neural models. As mentioned in Section 1, algebraic geometry provides richer invariants than differential and Riemannian geometry – such as the degree – and it is more suited for analyzing singular neuromanifolds. The importance of singularities for these models has led to the development of singular learning theory (Watanabe, 2009). The latter stems from information geometry, and naturally employs ideas from algebraic geometry. Neuroalgebraic geometry extends this direction, aiming to explore the full potential of a bridge between algebraic geometry and machine learning.
@@
Further examples of approaches bridging pure mathematics – in particular, algebra and geometry – with machine learning include:
  • geometric deep learning (Bronstein et al., 2021), concerned with symmetry properties of neural networks such as invariance and equivariance,
  • topological data analysis (Edelsbrunner & Harer, 2010), aiming at extracting topological features from data,
  • topological deep learning (Hajij et al., 2022), concerned with learning from data over non-Euclidean spaces, such as graphs and simplicial complexes,
  • categorical deep learning (Gavranović et al., 2024), leveraging on category theory for a compositional understanding of neural architectures.
@@ 3 Preliminaries
In this section, we review some basic concepts from machine learning and introduce the main objects of study of neuroalgebraic geometry, i.e., neuromanifolds.
@@
where $\mathcal{\ell}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a loss function. When $\mathcal{Y}$ is a Euclidean space or a probability simplex, respectively, common choices for $\ell$ are the quadratic loss $\ell(y,\hat{y})=\|y-\hat{y}\|^{2}$ or the cross-entropy loss $\ell(y,\hat{y})=-\sum_{j}y_{j}\log\hat{y}_{j}$. The goal of ERM is to find parameters that generalize to unseen examples. This is typically formulated by assuming that datapoints are drawn from some distribution, and that the empirical objective $L_{\mathcal{D}}$ approximates $\mathbb{E}_{x,y}[\ell(y,f_{w}(x))]$, referred to as generalization error. Often the data distribution is further assumed to take the form $x\sim\pi(x)$, $y=f^{*}(x)$, where $\pi$ is a distribution of inputs and $f^{*}$ is a deterministic function mapping inputs to outputs, interpreted as the ground-truth function underlying the given task.
In deep learning, sophisticated parametric models are constructed by composing simpler ones, referred to as modules or layers. The connectivity of such layers determines an architecture class – a family of models with similar design, e.g., a Multi-Layer Perceptron (MLP) is a sequential composition
@@ 3.1 Neuromanifolds
In other words, the neuromanifold is the image of the parametrization map, denoted by $\varphi\colon\mathcal{W}\rightarrow\mathcal{M}$, $w\mapsto f_{w}$. Describing neuromanifolds is challenging, as even simple architecture classes parametrize complex families of functions. Indeed, the parametrization map is typically non-linear and not one-to-one, resulting in different parameters that identify the same function. Therefore, the neuromanifold carries intrinsic geometric structure that differs drastically from the one of the parameter space $\mathcal{W}$. As we shall see, the geometry of $\mathcal{M}$ is related to several fundamental aspects in machine learning.
@@
The main class of examples of algebraic models is provided by neural networks with polynomial activation functions. Indeed, if $\sigma$ is a polynomial, then the expression in Equation 2 is polynomial in both $x$ and $w$. The study of MLPs with monomial activation function $\sigma(z)=z^{k}$ has been initiated by Kileel et al. (2019). For shallow networks – i.e., with $L=2$ layers – the corresponding neuromanifold coincides with the space of symmetric tensors of bounded (Waring) rank (Arjevani et al., 2025). For linear networks – i.e., with no activation function or, equivalently, with $k=1$ – the neuromanifold is, similarly, a space of matrices with bounded rank (Trager et al., 2020). This space is referred to as determinantal variety and it is well-understood – see Section A for a detailed discussion. For Convolutional Neural Networks (CNNs), the geometry of the neuromanifold differs drastically. In this case, up to rescaling the parameters, $\mathcal{M}$ is linearly birational – i.e., related by a linear map that is an isomorphism almost everywhere – to a more complex space known as the Segre-Veronese variety (Kohn et al., 2022, 2024; Shahverdi, 2024; Shahverdi et al., 2025a). A further example of an algebraic model is provided by linear (or ‘lightning’) attention mechanisms. These are non-normalized versions of the layers of a Transformer (Vaswani, 2017), which are popular nowadays across several applications. Linear attention mechanisms are cubical layers, and their neuromanifolds behave similarly to CNNs (Henry et al., 2025).
@@ 3.2 Learning on the Neuromanifold
The loss landscape is the graph of the loss $L_{\mathcal{D}}$ in parameter space. For algebraic loss functions $\mathcal{L}_{\mathcal{D}}$ (e.g., quadratic loss or Wasserstein distance for discrete distributions), the loss landscape is also a semi-algebraic variety, similar to the neuromanifold. For the cross-entropy loss or other loss functions with algebraic derivatives, at least the most important points of the loss landscape – namely, the critical points – depend algebraically on the data $\mathcal{D}$. Thus, in these settings, training is an algebraic optimization problem.
While the optimization in parameter space benefits from a Euclidean domain, enabling the use of algorithms such as gradient descent, the optimization problem in function space $\mathcal{V}$ is often more tractable from a mathematical perspective. In the algebraic setting, this problem is a polynomial program constrained to the neuromanifold. These programs have a well-developed theory, including, e.g., hierarchies of simplified relaxations (Lasserre, 2001). Moreover, optimization in function space can be often interpreted geometrically. For the quadratic loss, ERM consists in minimizing a potentially-degenerate quadratic form over $\mathcal{M}$, while the corresponding generalization error coincides with the squared $L^{2}$ distance between $f\in\mathcal{M}$ and the ground-truth function $f^{*}$. In both cases, learning amounts to a nearest-point problem over the neuromanifold w.r.t. the (squared) distance associated to the quadratic form. As anticipated in Section 1, this type of problems lies at the heart of metric algebraic geometry, whose tools, as we shall see, are central to neuroalgebraic geometry.
@@ 4.1 Dimension, Degree, and Covering Number
where $m$ and $d$ are the dimension and the degree of $\mathcal{M}$, respectively, and $C$ is a constant depending on the dimension of the ambient space $\mathcal{V}$ but independent of $m$, $d$, and $\varepsilon$. This type of bounds for covering numbers originated from the theory of Vitushkin variations (Yomdin & Comte, 2004) – higher-order analogues of the degree, capturing more meticolously the ‘twistedness’ of the variety.
@@ 4.2 Singularities
Interestingly, singularities of neuromanifolds frequently correspond to subnetworks of the original architecture, i.e., functions that can be represented by a smaller network from the same model class. This type of property has been observed for different architectures such as MLPs – with (Shahverdi et al., 2025b; Arjevani et al., 2025) and without (Trager et al., 2020) activation function – and for CNNs with polynomial activation functions (Shahverdi et al., 2025b). In these settings, singularities may be responsible for a form of ‘automatic model selection’, encouraging simpler models over more complex ones. This is also consistent with empirical observations in the literature, according to which the performance of a model can often be matched by sparse subnetworks obtained through pruning or distillation (Zhu et al., 2024) or with careful initialization (Frankle & Carbin, 2019).
@@ 4.3 Parameterization and Fibers
The fibers of a map $\varphi\colon\mathcal{W}\to\mathcal{M}$ describe domain points that are collapsed to the same output. Formally, the fiber of $f\in\mathcal{M}$ is $\varphi^{-1}(f)=\{w\in\mathcal{W}\ |\ \varphi(w)=f_{w}=f\}$. For an algebraic map $\varphi$, almost all its fibers resemble each other (over $\mathbb{C}$), and are collectively referred to as the generic fiber. Computing the fibers of a network’s parameterization formalizes the problem of identifiability in the context of machine learning, which is concerned with recovering the parameters of a model from the function they define.
@@
As an example, we now describe the fibers of MLPs. Other architectures have been considered, e.g. CNNs (Shahverdi et al., 2025b) and attention mechanisms (Henry et al., 2025). First, arbitrary MLPs exhibit combinatorial symmetries in their parameterization, by permuting the output of a layer and permuting back the input of the following one. Formally, given a permutation matrix $T$ for some layer $1\leq i<L$, the simultaneous transformations $W_{i}\mapsto TW_{i}$ and $W_{i+1}\mapsto W_{i+1}T^{-1}$ do not alter the function defined by the MLP. Moreover, for (positively) homogeneous activation functions $\sigma$ – e.g., $\sigma(x)=x^{r}$ or ReLU – a similar symmetry arises by rescaling the input and output of each neuron. Lastly, when $\sigma(x)=x$ is the identity, the same holds for arbitrary invertible matrices $T$. Intuitively, non-trivial activation functions ‘break’ symmetries, reducing the latter to neuron-wise transformations, i.e., permutations and possibly rescalings. It has been conjectured that such symmetries characterize the generic fibers of algebraic MLPs (Kileel et al., 2019; Kubjas et al., 2024), which has been recently proven for polynomial activations of high degree $r\gg 0$ (Finkel et al., 2024; Shahverdi et al., 2025b). In the non-algebraic case, the same result is well-known when $\sigma$ is the hyperbolic tangent (Fefferman et al., 1994), while for ReLU additional non-trivial symmetries arise (Grigsby et al., 2023).
@@ 4.4 Critical Points and Gradient Descent
In general, a parameter $w\in\mathrm{Crit}(\varphi)$ can be a critical point of the loss $L_{\mathcal{D}}$ even though the corresponding function $f_{w}\in\mathcal{M}$ is not a critical point for the loss functional $\mathcal{L}_{\mathcal{D}}$ in function space (see Section A for an example). These parameters are sometimes called spurious critical points (Trager et al., 2020). For some network architectures, spurious critical points will be typically avoided by gradient descent, while for others they are found with positive probability (Kohn et al., 2022). For instance, in the case of monomial activation, MLPs can have spurious critical points as local minima with positive probability (Kileel et al., 2019), while for CNNs they only correspond to the zero function $0\in\mathcal{M}$ (Shahverdi et al., 2025a). For attention networks, the characterization of spurious critical points is an open problem.
@@
Figure 4: Illustration of a curve with the corresponding data discriminant (dashed) and Voronoi cell of the singularity (gray). For points inside (resp. outside) the discriminant, the Euclidean distance from the curve has 2 (resp. 4) real critical points.
While critical points describe the equilibria of the gradient flow, its dynamical behavior can be described by its preserved quantities, referred to as dynamical invariants. These are relations in the parameters $w\in\mathcal{W}$ that stay constant along the curve that gradient flow traces in $\mathcal{W}$. Dynamical invariants have been discussed for a variety of network architectures (Kohn et al., 2022; Williams et al., 2019; Du et al., 2018), and the algebraic ones have been recently described in detail (Marcotte et al., 2024). Knowing the invariants not only enables to analyze the training dynamics, but can be exploited to design well-behaved parameter initializations, since gradient flow will avoid all parameters that do not satisfy the same invariants as the initial values (Arora et al., 2018b). For instance, a balanced initialization can lead to stable learning dynamics.